OpenStudy (anonymous):

The diameter, D , of a sphere is 15.8 . Calculate the sphere's volume, . Use the value 3.14 for Pi , and round your answer to the nearest tenth. (Do not round any intermediate computations.)

OpenStudy (saifoo.khan):

\[\frac43 \times \pi \times r^3\]

OpenStudy (anonymous):

lulz d = r/2 >< Not correct mister

OpenStudy (anonymous):

\[r=\frac{d}{2}\]

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

ummm ok